The perimeter of a rectangular cement patio is 36 meters. The area is 80 square meters. What are the dimensions of the patio? How do I solve this?

Answer by checkley79(3341)

P=2L+2W
A=LW
P=36
A=80
80=LW
L=80/W
36=2(80/W)+2W
36=160/W+2W
36=(160+2W*W)/W
36=(160+2W^2)/W
36W=160+2W^2
2W^2-36W+160=0
W^2-18W+80=0
(W-10)W-8)=0
W-10=0
W=10 ANS.
LW=80
L*10=80
L=80/10
L=8 ANS.
IF:
W=8 THEN:
L*8=80
L=80/8
L=10 ANS.
PROOF:
2*10+2*8=36
20+16=36
36=36